Use interest rate of 4.8894854% per year compounded monthly, find an appropriate value for A the makes the following two projects equivalent.

Project A Project B
Year Cash Flow Year Cash Flow

-4

120 0 35
-3 120 1 0
-2 100 2 A
-1 80 3 (1.07)*(A)
0 60 4 (1.07)^2 *(A)
1 40 5 (1.07)^3 *(A)
Answer #1

For the two projects to be equivalent the The value of both projects at time 0 should be equal

PV of A at year 0 = PV of B at Year 0

where r = 4.8894854%

120*(1+r)4 + 120*(1+r)3 + 100*(1+r)2 + 80*(1+r)1 + 60 + 40/(1+r) = 35*(1+r)4 + 0*(1+r)3 + A*(1+r)2 + 1.07A*(1+r)1 + 1.07^2 *A + 1.07^3 *A/(1+r)

575.7895 = 42.3638 + 4.535335 A

A = 117.62

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well